OpenStudy (anonymous):

A farmer has planted 75 acres of dryland wheat. He anticipates a yield of 60 bushels per acre and for each 6 bushels he hopes to be paid $ 8. Use dimensional analysis to find his total anticipated earnings.

OpenStudy (dumbcow):

6 bushels = $8 --> 1 bushel = $8/6 = $1.33 total bushels = 75*60 = 4500 total earnings = 4500*$1.33 = $6000